# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v3

@@ VI-A Observer variability
TABLE I: Comparison of test evaluations on two versions of the LIDC-IDRI dataset. Table adapted from [95].
@@
PU-Net [69], 2018 $0.310\pm 0.010$ $0.552\pm 0.000$ $0.320\pm 0.030$ $0.500\pm 0.030$
HPU-Net [74], 2019 $0.270\pm 0.010$ $0.530\pm 0.010$ $0.270\pm 0.010$ $0.530\pm 0.010$
PhiSeg [75], 2019 $0.262\pm 0.000$ $0.586\pm 0.000$
SSN [50], 2020 $0.259\pm 0.000$ $0.558\pm 0.000$
CAR [56], 2021 - - $0.264\pm 0.002$ $0.592\pm 0.005$
JProb. U-Net [135], 2022 - - $0.262\pm 0.000$ $0.585\pm 0.000$
PixelSeg [53], 2022 $0.243\pm 0.010$ $0.614\pm 0.000$ $0.260\pm 0.000$ $0.587\pm 0.010$
MoSE [81], 2022 $0.218\pm 0.001$ $0.624\pm 0.004$
AB [136], 2022 $0.213\pm 0.001$ $0.614\pm 0.001$
CIMD [96], 2023 $0.234\pm 0.005$ $0.587\pm 0.001$
CCDM [95], 2023 $0.212\pm 0.002$ $0.623\pm 0.002$ $0.239\pm 0.003$ $0.598\pm 0.001$
Explicitly modeling the annotator distribution has been explored with an MoE approach (Section V-C), using data with consistent image-annotation pairs, provided that annotators have an intrinsically associated expertise [137]. Nevertheless, relying on the model to infer ambiguity in the parameters by observing the data, can become quite burdensome. Hence, it can be much simpler to directly model the empirical stochasticity in the annotations, which has been extensively explored in modalities such as Lung nodule detection in 2D [69, 75, 74, 138, 139, 68, 60, 62, 73, 50, 59], as well as 3D [79, 140, 141], Brain Tumor [142, 50, 98, 94, 96], White Matter Hyperintensities [143], Pulmonary Tumour Growth [144], Prostate [145, 76] and vascular [146], street scene [147, 74, 95], aerial imaging [101], optic cup [99, 148], abdominal multi-organ [100] and nuclei microscopy [97] segmentation. The overwhelming majority of research for this particular application has been executed with variants of conditional VAEs [69, 60, 59, 73, 68, 62, 74, 75, 139, 144, 149, 142, 150, 141, 146, 79, 76, 145, 140]. More recently, the growing popularity of DDPMs is apparent in the field [95, 101, 98, 94, 96]. Also, GAN-based approaches have been employed [56].
VI-B Model introspection
The uncertainties obtained from probabilistic models can provide an insight into the reliability of a model when the correlation between uncertainty and model accuracy is strong. This relationship has been formalized by Mukhoti et al. [151] through two conditional likelihoods. Firstly, the accuracy given a certain prediction, $p(\text{A}|\text{C})$, and secondly, the uncertainty given an inaccurate prediction $p(\text{U}|\text{I})$. Given a threshold $u_{T}$ that distinguishes certain from uncertain pixels or patches, we can define pixels that are accurate and certain, accurate and uncertain, inaccurate and certain, inaccurate and uncertain, denoted by $u_{\text{ac}}$, $u_{\text{au}}$, $u_{\text{ic}}$, $u_{\text{iu}}$, respectively. Consequently, the authors combine $p(\text{A}|\text{C})=n_{\text{ac}}/(n_{\text{ac}}+n_{\text{au}})$ and $p(\text{U}|\text{I})=n_{\text{ic}}/(n_{\text{ic}}+n_{\text{iu}})$ to obtain the Patch Accuracy vs Patch Uncertainty (PAvPU) metric, defined as
This application has been applied in combination with MC dropout to model (video) street scenes [152, 153], remote sensing [154, 155], instance segmentation of various objects [156], point-cloud indoor scenes [157], brain [158, 159, 160, 17, 161, 162, 163], Multiple Sclerosis [164], cardiac [165, 166], heart ventricle [163], prostate [163], carotid artery [167], lumbosacral [19] MRI, and Optical Coherence Tomography [168, 169], skin imaging [170, 171], lung [171] and liver CT [172] and MRI  [173], and ultrasound [125, 174]. Concrete dropout has been applied to instance segmentation of C. elegans assays [175] and street-scenes [151]. To a lesser degree, ensembling [176, 125], Variational Inference both in 2D [177, 178] and 3D [179], M-heads (auxiliary networks) [14, 180, 181], and test-time augmentation [125] have also been used to quantify the uncertainty for quality assessment.
It can be noted that uncertainty is usually only obtained on pixel basis, while crucial information can be present in structural statistics. Hence, the Coefficient of Variation (CV) addresses this by measuring structural uncertainty through dividing the volume variance over the mean for all samples. Also, Roy et al. [162] propose to evaluate structural uncertainties by assuming predictions to be thresholded to binary masks with some function $t$, to determine the pair-wise average overlap between all respective samples as
$\overline{\mathrm{D}}=\mathbb{E}_{p_{\mathbf{Y}|\mathbf{X}^{*}}}\left[\,\text{Dice}(t(\mathbf{y}_{i}),t(\mathbf{y}_{j}))\,\right]\quad\forall i\neq j.$ (38)
Finally, including information on localized uncertainty to the training objective has shown to improve generalization capabilities [182, 183, 184, 185]. Note that using uncertainty to guide model training is closely related to Active Learning, which is discussed in Section VI-D.
VI-C Model generalization
As mentioned earlier in Section V-A, sampling new parameter permutations often improves segmentation performance, due to the model combining effect. For instance, dropout layers at the deeper decomposition levels of the SegNet [4] improves model performance [152]. Literature also reports improved performance with Concrete Dropout for semantic (instance) segmentation [151, 175]. This has been observed in multiple domains, including in-/outdoor scene understanding [152, 151, 186, 187, 181], brain tumor MRI [159, 188, 161, 177], Optical Coherence Tomography [168], low-dose Computed Tomography of lung nodules [189], colectoral polyps [190], cardiac MRI [18] and C. elegans roundworm microscopy images [175]. Ng et al. [18] have benchmarked multiple techniques for cardiac MRI segmentation and have found that ensembling results in the best performance improvement, while BBB is more robust to noise distortions.
Furthermore, the improved generalization from ensembling has also shown to produce more calibrated outputs [163]. In other work, orthogonality within and across convolutional filters of the ensemble is enforced through minimizing their cosine similarity, which reaped similar merits [49]. Nonetheless, individual models in conventional ensembles receive data in an unstructured manner. However, it is also possible to assign specific subsets of the data to particular models (so-called ‘experts’) in the ensemble [128], commonly referred to as a Mixture of Experts (MoE). While MoE resembles ensembling in many ways, the approach additionally relies on a learnable gate that inherits the decision-taking logic. Pavlitskaya et al. [181] show merits of such approach for urban outdoor scene segmentation. For optic cup segmentation, Ji et al. [137] also condition on a normalized expertness vector, where each element corresponds to the weight given to a particular expert, and is inserted at the deepest layer of a U-Net. Gao et al. [81] have introduced a Mixture of Stochastic Experts (MoSe), which can be regarded as a stochastic adaptation of the MoE approach. Nevertheless, the methodology addresses aleatoric uncertainty quantification and has hence been discussed in Section IV-B2.
VI-D Active Learning
The field of active learning [191, 192] aims to reduce the costly annotation procedure by careful selection of unlabeled training samples. A wide range of methodologies exist, but since the nature of this problem involves identifying and reducing model ignorance, the quantification of epistemic uncertainty is most appropriate. In terms of metrics, the expected increase in posterior entropy, defined by
$H[\,\bm{\theta}|\mathcal{D}\,]-\mathbb{E}_{p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})}[\,H[\,\bm{\theta}|\mathbf{y},\mathbf{x}^{*},\mathcal{D}\,]\,],$ (39)
can provide a notion to describe information gain, and thus the decrease of uncertainty for specific datapoints. Notably, minimizing the expected increase in posterior entropy is equivalent to maximizing the mutual information between the data and model parameters. This enables reformulation in terms of the output space, rather than the complex parameter space and is known as Bayesian Active Learning by Disagreement (BALD) [193], denoted as
$I(\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D})$ $=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-H[\mathbf{y}|\bm{\theta},\mathbf{x}^{*},\mathcal{D}]$
$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,].$ (41)
Active Learning through uncertainty quantification has been been beneficial for medical [194, 195, 196, 197, 198, 199, 200, 201], multi-view [202], remote sensing [203], street-views [204] and 3D point-cloud data [205]. These works mostly utilize MC Dropout to obtain a notion of uncertainty, but explicit VI together with BALD [185] and novel temporal-ensembling methods [206] have also been used to reduce the required human labeling. Some methods extend beyond pixel-level uncertainty, incorporating boundary information in the uncertainty quantification process [204, 200]. For example, Kasarla et al. [204] make use of both image and pixel-level uncertainty. Furthermore, pixels are weighed by their closeness to edges, as boundary cases are more likely to be uncertain. Similarly, Ma et al. [200] combine target and boundary-based uncertainty sampling, thereby ensuring diversity, effective and balanced utilization of the available information.
VII Discussion
This section highlights the main trends and discussion points in the field. The challenges and limitations related to the methods (Section VII-A) and downstream applications (Section VII-B) are discussed. Based on these observations, recommendations for future work will be provided in Section VII-C.
VII-A Discussion on methods
The core challenges with modeling segmentation distributions entails obtaining calibrated uncertainty estimates and secondly, designing spatial coherence between pixels of the samples. While the former can be achieved with conventional CNN models, the latter requires modeling correlation across pixels of the segmentation mask. As discussed in earlier sections, the desired nature of the uncertainty estimates governs the required methodology. Similarly, our discussion will follow a comparable format as the preceding sections, i.e. the aleatoric and epistemic approaches will initially be treated in a separate manner.
VII-A1 Aleatoric methods
The literature overview indicates that three distinct routes can be taken to model the correlation. The first method is to model the correlation directly in the pixel space. The second method entails using latent-variable modeling, where the correlation is encoded in the latent space. Finally, TTA is a more straightforward and practical approach to obtain uncertainty and requires the least modifications to existing models. We will discuss the strengths and weaknesses of each of those approaches. A summary of this is presented in Table II.
@@ VII-A1 Aleatoric methods
TABLE II: Comparison between models that quantify aleatoric uncertainty.
@@
TTA Model agnostic, no additional training parameters Implicit likelihoods [103, 104, 105, 106]
@@
SSN Model agnostic, explicit likelihoods, fast sampling Unstable training [50]
@@
PixelCNN Explicit, exact likelihoods Sequential sampling, memory intensive [53]
@@
GAN Fast sampling, flexible Unstable training, poorly defined objective, implicit likelihoods [56, 57]
VAE Fast sampling, flexible, interpretable latent space, ELBO Mode/posterior collapse, amortization gap [69, 74, 75, 59, 60, 73, 62, 61, 141, 89, 81, 135, 68]

### v4

@@ 7.3 Standardization
Table 6: Comparison of test evaluations on two versions of the LIDC-IDRI dataset. Table extended from (Zbinden et al., 2023). Indicates evaluation results not obtained from the original work.
@@
PU-Net (Kohl et al., 2019b) $0.310\pm 0.010^{*}$ $0.552\pm 0.000^{*}$ $0.320\pm 0.030^{*}$ $0.500\pm 0.030^{*}$
HPU-Net (Kohl et al., 2019a) $0.270\pm 0.010^{*}$ $0.530\pm 0.010^{*}$ $0.270\pm 0.010^{*}$ $0.530\pm 0.010^{*}$
PhiSeg (Baumgartner et al., 2019) $0.262\pm 0.000^{*}$ $0.586\pm 0.000^{*}$
SSN (Monteiro et al., 2020) $0.259\pm 0.000^{*}$ $0.558\pm 0.000^{*}$
CAR (Kassapis et al., 2021) - - $0.264\pm 0.002^{\phantom{*}}$ $0.592\pm 0.005^{\phantom{*}}$
JProb. U-Net (Zhang et al., 2022d) - - $0.262\pm 0.000^{\phantom{*}}$ $0.585\pm 0.000^{\phantom{*}}$
PixelSeg (Zhang et al., 2022c) $0.243\pm 0.010^{\phantom{*}}$ $0.614\pm 0.000^{\phantom{*}}$ $0.260\pm 0.000^{\phantom{*}}$ $0.587\pm 0.010^{\phantom{*}}$
MoSE (Gao et al., 2022) $0.218\pm 0.001^{\phantom{*}}$ $0.624\pm 0.004^{\phantom{*}}$
AB (Chen et al., 2022) $0.213\pm 0.001^{*}$ $0.614\pm 0.001^{*}$
CIMD (Rahman et al., 2023) $0.234\pm 0.005^{*}$ $0.587\pm 0.001^{*}$
CCDM (Zbinden et al., 2023) $0.212\pm 0.002^{\phantom{*}}$ $0.623\pm 0.002^{\phantom{*}}$ $0.239\pm 0.003^{\phantom{*}}$ $0.598\pm 0.001^{\phantom{*}}$
BerDiff (Chen et al., 2023) - - $0.238\pm 0.010^{\phantom{*}}$ $0.596\pm 0.000^{\phantom{*}}$
MedSegDiff (Wu et al., 2023b) - - $0.420\pm 0.030^{*}$ $0.413\pm 0.030^{*}$
SPU-Net (Valiuddin et al., 2024b) - - $0.327\pm 0.003^{\phantom{*}}$ $0.560\pm 0.005^{\phantom{*}}$
7.4 Future of Active Learning
Among the applications considered, Active Learning emerges as a promising approach to uncertainty quantification, showing true potential of uncertainty quantification by potentially reducing annotation costs while preserving or even improving model generalization. Active Learning directly addresses the labeling bottleneck through uncertainty-driven selection, aligning with model introspection and fostering human-AI collaboration in specialized domains. Furthermore, when combined with federated learning, this field can accelerate privacy-centric collaboration by enabling efficient, human-in-the-loop improvement of safety-critical models across decentralized data silos. However, key limitations remain and require further research.
Model sensitivity and baselines
While deep learning-based segmentation typically involves high-dimensional data and large datasets, active learning requires a well-generalized model trained on limited data. This mismatch renders performance highly sensitive to budget constraints, model architecture, hyperparameters, and regularization Mittal et al. (2019); Kirsch (2024); Munjal et al. (2022); Kirsch et al. (2019); Gaillochet et al. (2023). Furthermore, random sampling, particularly when combined with MC Dropout, often proves to be a surprisingly strong baseline that is difficult to outperform Kahl et al. (2024); Mittal et al. (2019); Lüth et al. (2023); Burmeister et al. (2022); Siddiqui et al. (2020); Kim et al. (2021); Sinha et al. (2019); Nath et al. (2020), especially with highly imbalanced data Ma et al. (2024a). This observation can partly be attributed to the fact that MC Dropout likely fails to capture true Bayesian uncertainty (see Section 7.7), inviting exploration of other techniques.
Sample informativeness
Besides uncertainty, the diversity of selected samples is equally critical. For instance, a set of highly uncertain samples may share similar semantic content, leading the model to overfit a narrow and homogeneous region of the data distribution during early training. Hence, sample selection should be based on both uncertainty and diversity (Ozdemir et al., 2021; Wu et al., 2021; Jensen et al., 2019; Burmeister et al., 2022). Ozdemir et al. (2021) leverage the latent space of a VAE to select representative samples, whereas other studies use feature vectors from prediction models (Wu et al., 2021; Burmeister et al., 2022). Owing to their significantly greater representational capacity, modern generative models, as discussed in Section 7.6, merit experimental investigation together with segmentation specific conditioning. Furthermore, the sample diversity offered by such models can be effectively combined with adversarial training (Mahapatra et al., 2018; Sinha et al., 2019).
Uncertainty calibration and aggregation
As discussed in Section 4.1, using the predictive entropy of SoftMax-based uncertainty requires proper calibration. Unfortunately, this is rarely verified in practice Kasarla et al. (2019); García Rodríguez et al. (2020); Xie et al. (2022); Wu et al. (2021); Burmeister et al. (2022); Gaillochet et al. (2023). Furhtermore, most works rely on heuristics to aggregate the pixel-level information such as closeness to edges, boundaries or the use of super-pixels Gorriz et al. (2017); Kasarla et al. (2019); Ma et al. (2024b); Gaillochet et al. (2023); Wu et al. (2021); Kasarla et al. (2019); Bengar et al. (2021). Moreover, the mutual information is often quantified by plain summation (i.e. assuming pixel independence) Shen et al. (2021); Ma et al. (2024b). We hypothesize that these ad-hoc aggregation techniques, as mentioned in Section 7.2, are a key limiting factor to uncertainty-based Active Learning.
The Cold-Start Problem
Finally, the “cold start” problem Nath et al. (2022); Houlsby et al. (2014); Yuan et al. (2020a), the notion that uncertainty methods work poor on the initial samples, is rarely addressed in most work. Most approaches assume a well-trained model is already available, yet methods like MC Dropout require warm-up epochs to yield meaningful uncertainty. Addressing this gap, post-hoc techniques such as that of Sourati et al. (2018), which enable uncertainty estimation from pretrained deterministic networks, represent a valuable and underexplored direction for early-stage sample selection.
7.5 Segmentation backbone
Due to emphasis on the uncertainty methods, the backbone feature extractor is often an overlooked element in uncertainty modeling. For example, CNN-based encoder-decoder models such as the U-Net, remain the preferred backbone architecture (Eisenmann et al., 2023). Nevertheless, Vision Transformers (ViTs) (Vaswani et al., 2017) such as SegFormer, Swin U-Net, Mask2Former already outperform CNNs in general segmentation problems (Xie et al., 2021; Zhang & Yang, 2021; Hümmer et al., 2023). With the exception of experiments in Kahl et al. (2024) (combinding SSN (Monteiro et al., 2020) with HRNet (Yuan et al., 2020b)), all employed models use a CNN backbone. The limited adoption of ViTs likely stems from CNNs’ advantageous inductive biases, whereas Transformers typically require extensive pretraining on large datasets (Caron et al., 2021). Therefore, we recommend integrating the discussed methods into similar architectures to achieve both strong classification performance and effective ambiguity modeling. Also, CNNs have benefited from the recent developments in Transformers. For instance, “ConvNext" takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains (Liu et al., 2022c).
7.6 Modeling feature distributions
In many instances, latent-variable models are employed to model the observer variability. To aid with appropriate model selection, we will discuss the strengths and weaknesses of each of those approaches. A summary of this is presented in Table 7. Note that while classified PixelCNNs and SSNs are classified as ‘pixel-level’ methods, both can be phrased as a latent-variable model as well.
ELBO suboptimality
Known for its flexibility to various datasets, fast sampling time and the interpretable latent space, VAEs seem to be the most popular choice for aleatoric uncertainty quantification. Nonetheless, the shortcomings of VAEs are well known. For example, such models suffer from inference suboptimality related to ELBO optimization Cremer (2018); Zhao et al. (2017) and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of model collapse Valiuddin et al. (2024a; b); Qiu & Lui (2020). This has been hypothesized to be caused by excessively strong decoders Chen et al. (2016) and is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective Kohl et al. (2019a), residual connections Kohl et al. (2019a); Baumgartner et al. (2019), or deep supervision Baumgartner et al. (2019) are required for generalization. A unique benefit of this approach is the ability to semantically interpret the latent space with, for example, interpolation between annotation styles or the exploration of low-likelihood regions.
Limitation of sequantial modeling
The adoption of sequential sampling models DDPMs or PixelSeg is rather limited compared to the VAE-based models. This is unexpected, as both models outperforms the latter. This is likly due to their tedious sequential inference procedure Song et al. (2020); Zheng et al. (2023); Meng et al. (2023); Zhang et al. (2022c), a crucial limitation that is exacerbated in supervised settings, which often require validation through sampling on a separate data split. Despite these shortcomings, the strengths of DDPMs should not be overlooked. Their iterative sampling allows for highly flexible modeling and preserves high-frequency details often lost in latent-variable approaches like VAEs, which can produce blurry reconstructions. Moreover, with existing work on faster sampling methods such as DDIMs (Song et al., 2020) and Flow Matching Lipman et al. (2022), these models warrant further investigation.
Discrete vs. continuous
It can be noted in Table 6 that the best-performing DDPM models are discrete in nature Chen et al. (2023); Zbinden et al. (2023). While it is debatable whether shifting to categorical distributions is required for complex image generation Chen et al. (2022), this observation implies the need for further investigation of the merits of categorical distributions in segmentation settings. Visually, this can be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13).
(a) LIDC-IDRI Armato (2011)
(b) CityScapes Cordts et al. (2016a).
Figure 13: Continuous vs. categorical forward diffusion process with cosine noise scheduling Nichol & Dhariwal (2021). Note the use of categorical diffusion that results in a more gradual transition for the multi-class case.
Others
SSNs are a simple, fast, and model-agnostic approach for modeling uncertainty. However, they can suffer from training instability due to the requirement that the covariance matrix be invertible. To address this, the authors propose masking out the background to avoid exploding variances and using uncorrelated Gaussians when the covariance matrix is singular. Despite these fixes, SSNs scale poorly with data, and a potential solution is to shift the stochasticity deeper into the network by introducing a latent prior, resembling VAE-based models but with a more complex latent space. CAR, on the other hand, is a less commonly used method, likely due to the well-known instability of GAN-based models Arjovsky & Bottou (2017). GANs often require additional heuristic terms in the loss function to ensure stable training. CAR is no exception to this, as its objective function combines four separate losses.
Recommendation
We advocate for a deeper contextualization of contemporary generative modeling research within probabilistic segmentation models. Generative modeling is a rapidly evolving field and any such model can be modified for segmentation through intricate conditioning. For example, DDPMs were initially proposed for general image generation (Ho et al., 2020), but was quickly adopted for stochastic segmentation. Furthermore, flow-based models also enjoy many benefits yet to be explored in this field, such as exact likelihood modeling and faster sampling Lipman et al. (2022). In terms of model selection, we have noticed SSN being the most practical solution across literature due to being model agnostic and receiving good benchmarking scores (Kahl et al., 2024; Monteiro et al., 2020; Ng et al., 2022)
@@ 7.6 Modeling feature distributions
Table 7: Comparison between generative models that model aleatoric uncertainty.
@@
SSN Model agnostic, explicit likelihoods, fast sampling possibly unstable, memory intensive (Monteiro et al., 2020; Kahl et al., 2024; Zepf et al., 2023a; Philps et al., 2024; Kahl et al., 2024)
@@
PixelCNN Explicit, exact likelihoods Sequential sampling, memory intensive Zhang et al. (2022c)
@@
GAN Fast sampling, flexible Unstable training, poorly defined objective, implicit likelihoods (Kassapis et al., 2021; Isola et al., 2017)
VAE Fast sampling, flexible, interpretable latent space, ELBO Mode/posterior collapse, amortization gap (Kohl et al., 2019b; Selvan et al., 2020; Valiuddin et al., 2021; 2024a; Qiu & Lui, 2020; Bhat et al., 2022a; 2023; Valiuddin et al., 2024b; Zhang et al., 2022b; Kohl et al., 2019a; Baumgartner et al., 2019; Hu et al., 2023; Liu et al., 2022b; Fischer et al., 2023; Savadikar et al., 2021; Rafael-Palou et al., 2021; Viviers et al., 2023b; a; Chotzoglou & Kainz, 2019; Saha et al., 2021a; b; Long et al., 2021a; Schmidt et al., 2023; Savadikar et al., 2021; Rafael-Palou et al., 2021; Zepf et al., 2024; 2023a; Hu et al., 2022)